OpenStudy (anonymous):

what is the lcm of 35c^2b^4, 5c^6b^8, 21c^2b^3

OpenStudy (anonymous):

35c^2b^4 = 5 * 7 * c * c * b * b * b * b 5c^6b^8 = 5 * c * c * c * c * c * c * b * b * b * b * b * b * b * b 21c^2b^3 = 3 * 7 * c * c * b * b * b LCM = 3 * 5 * 7 * c^6b^8 LCM = 105c^6b^8 (always use the highest exponent on each variable) (multiply all numbers together if two or more terms have the same factor... ex. 5 and 7 up above.. you only count it once)